OpenStudy (anonymous):

A pension fund owns 2,400 fewer shares in mutual stock funds than mutual bond funds. Currently, the stock funds sell for $14 per share, and the bond funds sell for $17 per share. How many shares of each does the pension fund own if the value of the securities is $146,200?

OpenStudy (anonymous):

Need to find amount of: stock funds shares bond fundsshares

OpenStudy (paxpolaris):

:) ... So let's choose some variables and write some equations.

OpenStudy (anonymous):

ok

OpenStudy (paxpolaris):

can we use two variables? S is for stocks, B is for bonds.

OpenStudy (anonymous):

ok

OpenStudy (paxpolaris):

A pension fund owns 2,400 fewer Stocks than Bonds. ... can you translate this into an equation using S and B

OpenStudy (anonymous):

B-2400

OpenStudy (paxpolaris):

right the number of Stocks is equal to B-2400 \[\large S= B-2400\]

OpenStudy (paxpolaris):

now for the second equation: Stocks have a value of $14 each, Bonds have value $17 each.... And the total value is $146,200.

OpenStudy (anonymous):

Ok so am I basically dividing 14 and 17 into it

OpenStudy (paxpolaris):

? we need to write another equation using S and B.

OpenStudy (anonymous):

ok

OpenStudy (paxpolaris):

1 Stock have a value of \(14\) so, S stocks will have value of \(14S\)

OpenStudy (anonymous):

ok

OpenStudy (paxpolaris):

same way, 1 Bond has a value of \(17\) so, B stocks will have value of \(B\times17=17B\)

OpenStudy (anonymous):

ok

OpenStudy (paxpolaris):

add both \(14S\) and \(17B\) up .... and we should get the total value of \(146200\)

OpenStudy (paxpolaris):

so the 2nd equation is :\[14S+17B=146200\]

OpenStudy (anonymous):

ok that makes sense

OpenStudy (paxpolaris):

so we have: \[S=B-2400\] and, \[14S+17B=146200\] can you solve for S and B from here?

OpenStudy (anonymous):

not quite

OpenStudy (paxpolaris):

well since we know that \(S\) is the same number as \(B-2400\) from the first equation. in the 2nd equation where we see \(S\) we can replace it with \((B-2400)\)

OpenStudy (paxpolaris):

\(14S+17B=146200\) will becomes: \[\large \implies14(B-2400)+17B=146200\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

how about the other one?

OpenStudy (paxpolaris):

so what do you have for the number of mutual bond fund shares

OpenStudy (anonymous):

5800

OpenStudy (paxpolaris):

\[\huge \color{green}\checkmark\] so to find the number of stocks you just have to go back to the first sentence.

OpenStudy (anonymous):

5800-2400

OpenStudy (anonymous):

3400

OpenStudy (paxpolaris):

A pension fund owns 2,400 fewer Stocks than Bonds. \[S=B-2400\\=5800-2400\\=3400\] great job :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you were very helpful!

OpenStudy (paxpolaris):

yw